OpenStudy (anonymous):

A soup can has a height of 4.25 inches and a diameter of 3.25 inches. Which of the following is the closest to the volume of the soup can? 46 cubic inches 141 cubic inches 22 cubic inches 35 cubic inches

OpenStudy (anonymous):

will medal n fan

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@coolman500

OpenStudy (anonymous):

r u good at math

OpenStudy (anonymous):

Well, we know it is a “soup can”, so we are finding the volume of a cylinder. The formula for volume of a cylinder is \[\pi(r ^{2}h)\] So substitute the equation with your given values.

OpenStudy (anonymous):

You are given with a diameter, but we need to find the radius, and since the radius is half of the diameter, divide 3.25 by 2: 3.25/2=1.625 So 1.625 is your radius, now you can find the volume

OpenStudy (anonymous):

\[\pi(1.625^{2})(4.25)\] = pi times (2.640625)(4.25), which is about 11.22265625 pi. Multiply 11.22265625 by pi and you get 35.2570144288, or about 35.3, or 35 cubic inches.

OpenStudy (anonymous):

So, your final answer is D. 35 cubic inches. @slaay.y

OpenStudy (anonymous):

..Happy studying! :)